OpenStudy (anonymous):

Harris has two pipes to fill aquarium. Pipe A can fill the aquarium at the rate of 4 gallons per minutes. Pipe B can fill the aquarium at the rate of 840 liters per our. [1 gallon=3.79 liters] Part A: Which pipe will fill the aquarium faster? Justify your answer by comparing the rates after converting to the same units.

OpenStudy (anonymous):

will give medal and fan

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

PIPE A will fill the Aquarium Faster 4 * 3.79 = 15.16 15.16 * 60 = 909.6 909.6 > 800

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

It is PIPE A

OpenStudy (anonymous):

I have one more question? Can you help?

OpenStudy (anonymous):

Thanks For the Medal!

OpenStudy (anonymous):

Is that all?

OpenStudy (acxbox22):

Pipe A can fill 240 gallons per hour and pipe B can fill 840 liters in an hour 240x3.79=909.6 Pipe A can fill it faster because 909.6 is greater than 840.

OpenStudy (anonymous):

I have one more question. Can you help?

OpenStudy (anonymous):

What's Up?

OpenStudy (anonymous):

Part B: If the total volume of the aquarium was 500 gallons and you checked on how full it was every 0.25 hours, describe appropriate scales on the x-axis and y-axis to graph the rate at which the faster pipe fills the aquarium.

OpenStudy (anonymous):

500 gallons * 3.79 = 1895 1895 / 909.6 = 2.08 hours

OpenStudy (anonymous):

Just a sec I'll Attach A file.

OpenStudy (anonymous):

Here...

OpenStudy (anonymous):

Okay thanks. I gave you a medal and fan

OpenStudy (anonymous):

I hope This Helped :)

OpenStudy (anonymous):

<oo> \____/

OpenStudy (anonymous):

Bye Now!

OpenStudy (anonymous):

Can you explain the describe it for me. I still don't understand

OpenStudy (anonymous):

Uhhh Ok, Y = Liters X = Hours The Line is how many liters( Y )fill up the tank in ( X ) Hours.

OpenStudy (anonymous):

Sorry, i meant like can you help me describe appropriate scales on the x-axis thing.

OpenStudy (anonymous):

Oh OK, So If you look Under the X axis you should see numbers in this Order 0 0.5 1 1.5 2 Those should be he scale, beacause the intervals are "You check the tank every 0.25 hours" You can see the 0.25, 0.5, 0.75, 1.0, 1.25, 1.5, 1.75, and 2.0.

OpenStudy (anonymous):

ok, is that the description?

OpenStudy (anonymous):

It is a Description, and should work out if you awnser that, there is no garentee, but I think your teacher will except that awnser. I hope it helps <o o> ~

OpenStudy (anonymous):

ok, thanks so much

OpenStudy (anonymous):

Cool, Bye!

OpenStudy (anonymous):

I understand it now